OpenStudy (anonymous):

Use properties of logarithms to expand the loarithmic expressions as much as possile. log4 7-2 1) 7log4 4 2) 4log4 7 3) -4log4 7 4) -16log 7

OpenStudy (anonymous):

do you mean log_4(7^(-2)) ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

properties of log log_4(7^(-2)) rule: log(x^(y) = ylog(x) thus, log_4(7^(-2)) = -2log_4(7) I dont see the answer there

OpenStudy (anonymous):

the answer is -2log_4 7

OpenStudy (anonymous):

I dont think this question provides any correct solutions to this problem if anyone disagrees please reply

OpenStudy (anonymous):

If it is an online assignment i would send a message to the prof or select number 3 as it is the closest to the right answer

OpenStudy (anonymous):

there is no ryt answer in your choices